OpenStudy (anonymous):

14. For questions 14–19, what are the solutions of the equation? z2 – 10z + 24 = 0 (1 point) –4, –6 4, –6 4, 6 –4, 6 15. 2z2 – 4z + 2 = 0 (1 point) 2, –1 , 1 , –1 2, 1

OpenStudy (luigi0210):

For 14 just factor it. (z-6)(z-4) because both add up to -10 and multiply to +24. Now just set each equal to 0 and solve. z-6=0 z-4=0

OpenStudy (anonymous):

um C ?

OpenStudy (luigi0210):

It would help if they were labeled but yes. Another method would be to use the quadratic formula, or graph it and find where it intersects the x-axis.

OpenStudy (anonymous):

thanks(: do you know 15 ?

OpenStudy (luigi0210):

Same concept it just involves an extra step. So first we factor out a 2: \[2(z^2-2z+1)\] now set it equal to zero and divide by the 2, and since anything divided by 0 is 0 it would change anything: \[z^2-2z+1=0\] Now just factor and solve: \[(z-1)(z-1)\] And again both add up to a -2 and multiply to a +1.

OpenStudy (luigi0210):

*wouldn't change anything

OpenStudy (anonymous):

Gotcha so its either 1,1 or 1,-1

OpenStudy (luigi0210):

It's 1, 1 because there are no z+1 in the factored equation

OpenStudy (anonymous):

ok gotcha(: Can you help me with a few more please ?

OpenStudy (luigi0210):

Sure.

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

6. c2 – 10c = 0 a. 0, 10 b .0, over 10 c. 0, –10 d. 1, –10 15 = 8x2 – 14x a. -4, 3/8 b. -2/5, 4/3 c. -3, 5/8 d. -3/4, 5/2

OpenStudy (anonymous):

6x2 – 17x + 13 = 20x2 – 32 a. 5/3, 9/2 b. -5/2, 9/7 c. -5/2, -9/7 d. -5/3, 9/2 20x2 + 7x – 3 = 0 a. -3/5, 1/4 b. 3/5, 1/4 c. -3/5, -1/4 d. 3.5, -1/4

OpenStudy (luigi0210):

Well let's start with 6. What can we do first, any ideas?

OpenStudy (luigi0210):

It's like the previous problem except it's just a variable.

OpenStudy (anonymous):

ummm

OpenStudy (luigi0210):

:/ Factor out a c: \[c(c-10)=0\] Now separate them and solve: \[c=0\] \[c-10=0\]

OpenStudy (anonymous):

im sorry lol i suck at alg ! :( okay hold on

OpenStudy (anonymous):

im going to go with B ?

OpenStudy (luigi0210):

it's a :/

OpenStudy (anonymous):

hey hey i was close lol (:

OpenStudy (luigi0210):

Yup Okay next one: 15=8x^2-14

OpenStudy (anonymous):

worked it out of paper i think its C or D

OpenStudy (luigi0210):

How did you start?

OpenStudy (anonymous):

started with this 8x^2-14

OpenStudy (luigi0210):

You always start by moving everything to one side so it can equal 0.

OpenStudy (luigi0210):

so it should be \[8x^2-14x-15\]

OpenStudy (anonymous):

yes ^

OpenStudy (luigi0210):

Now, since this is almost impossible to factor mentally, we have use the quadratic formula \[\frac{ -b+/-\sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

Gotha , what do we start with first ?

OpenStudy (luigi0210):

We just plug in numbers and start solving it. In this equation the a=8, b=-14 and c =-15

OpenStudy (anonymous):

im confuesed

OpenStudy (luigi0210):

it would look like this: \[\frac{ 14+\sqrt{-14^2-4(8)(-15)} }{ 2(8) } ; \frac{ 14-\sqrt{-14^2-4(8)(-15)} }{ 2(8) }\]

OpenStudy (luigi0210):

What are you confused about?

OpenStudy (anonymous):

i got it look

OpenStudy (anonymous):

15-15=8x^2-14x-15 0=8x^2-14x-15 0=(4x+3)(2x-5) 4x+3-3=0-3 4x/4=-3/4 x=-3/4

OpenStudy (anonymous):

2x-5+5=0+5 2x/2=5/2 x=5/2

OpenStudy (luigi0210):

Oh, haha I guess you did. Sorry I made it a bit harder than it needed to be. Just forget what I said earlier :P

OpenStudy (anonymous):

lol its fine , skip all those problems bc i got them.

OpenStudy (anonymous):

What are the solutions of the equation 2x2 = 2? Use a graph of the related function

OpenStudy (luigi0210):

Well put it in standard form.

OpenStudy (luigi0210):

Aren't equation 3 and 1 the same?

OpenStudy (anonymous):

sorry ,

OpenStudy (anonymous):

2x2 = 2 4 = 2

OpenStudy (anonymous):

wait wait oppss i did it wrong!

OpenStudy (anonymous):

2x2=8 2x2−8=0 y=2x2−8

OpenStudy (luigi0210):

Where did that 8 come from..?

OpenStudy (anonymous):

bc 2 & 2 ? or the 4 should b in the 8's place

OpenStudy (luigi0210):

:/

OpenStudy (anonymous):

lol dont gimme that face!

OpenStudy (luigi0210):

Ha, sorry I just that the original equation was 2x^2=2

OpenStudy (anonymous):

and you get 1

OpenStudy (luigi0210):

well rad 1 but yea

OpenStudy (anonymous):

-1 , 1 C

OpenStudy (luigi0210):

equation ?

OpenStudy (luigi0210):

3?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

Yup

OpenStudy (anonymous):

Graph the function. Identify the vertex and axis of symmetry. f(x) = 4x2 –4x – 2

OpenStudy (luigi0210):

The vertex is -b/2a